Question 1

Under the tax rate of 40%, an unlevered firm Pacific Company’s WACC is currently 15 percent. The company can borrow at 6 percent. 

a.  What is Pacific Company’s cost of equity? 

b.  If the firm converts to 30 percent debt, what will its cost of equity be? 

c.  If the firm converts to 55 percent debt, what will its cost of equity be? 

d.  What is Pacific Company’s WACC in part (b)? In part (c)?  

 

 

Q 2

Smart Toys, currently has no debt, expects an EBIT of $45,000 every year forever. Its cost of equity is 16 percent. The corporate tax rate is 30 percent. The company can borrow at 8 percent.

a.​What is the current value of the company? 

b.​What will the value of the firm be if the company takes on debt equal to 20 percent of its unlevered value? What if it takes on debt equal to 60 percent of its unlevered value? 

c. ​What will the value of the firm be if the company takes on debt equal to 55 percent of its levered value? What if the company takes on debt equal to 40 percent of its levered value?  



1
Expert's answer
2022-02-24T11:36:03-0500

Question 1

WACC= "\frac{E}{V}\times K_e + \frac{D}{V}\times K_d(1-tax)"

Where E= Value of equity 

D=Value of debt

"K_e" = Cost of equity 

"K_d" = Cost of debt 

V= Total value of firm = Debt+Equity

By substituting given values in above formula we get,

(a) The firm is unlevered which means there is no debt in its capital structure so debt will be 0.

"0.14=1\times K_e"

"K_e" = 0.14 or 14%

 

(b)"K_e=K_o+(K_o-K_d)\frac{D}{E}\times(1-Tax)"

="0.14+(0.14-0.06)\frac{0.3}{0.7}(1-0.40)"

=0.14+0.0206=0.1606 or 16.06%

 

(c)  D="55\%" so E=1-D="45\%"

"K_e" =0.14+(0.14-0.06)0.55/0.45(1-0.4)

=0.14+0.0586=0.1986 or 19.86%

 

(d) 

WACC when D=0.30

WACC="0.1606\times 0.70+0.30\times0.06\times(1-0.4)"

WACC=0.11242+0.0108=0.1232 or 12.32%

 

WACC in  (c) Where D=0.55

WACC="0.1986\times0.45+0.55\times0.06\times(1-0.4)"

WACC=0.08937+0.0198=0.10917 or 10.92%


Question 2


Step-by-step explanation

a. What is the current value of the company? 

Value="K_e"


EBIT(1−Tax)

="\frac{45,000(1−0.3)}{0.16}"

=196,875

 

b-1. What will the value of the firm be if the company takes on debt equal to 20 percent of its unlevered value?

 

Debt=20%"\times" 196,875=39,375

 

Value of Levered Firm=Value of unlevered firm+(debt"\times" Tax rate)

=196,875+(39,375"\times" 30%)

=196,875+11,812.5

=208,687.5

 

b-2 What if it takes on debt equal to 60 percent of its unlevered value?

 

Debt=60"\%\times" 196,875=118,125

 

 

Value of Levered Firm=Value of unlevered firm+(debt∗Tax rate)

=196,875+(118,125"\times" 30"\%" )

=196,875+35,437.5

=232,312.5

 

c-1. What will the value of the firm be if the company takes on debt equal to 55 percent of its levered value?

Value of Levered Firm=Value of unlevered firm+(debt"\times" Tax rate)

VL=196,875+(0.55VL"\times" 30"\%" )

VL=196,875+0.165VL

VL−0.165VL=196,875

0.835VL=196,875

VL=235,778.44

 

c-2 What if the company takes on debt equal to 40 percent of its levered value? 

Value of Levered Firm=Value of unlevered firm+(debt"\times" Tax rate)

VL=196,875+(0.40VL"\times" 30%)

VL=196,875+0.12VL

VL−0.12VL=196,875

0.88VL=196,875

VL=223,721.59